### v2

@@ 4.2 Convection-Diffusion problem
where $u_{i}$ represents the numerical approximation to $u(x_{i})$. Let $M$ be the tridiagonal matrix obtained from the above finite-difference discretization of $L$. We take $\bm{b}$ to be the vector with all entries equal to one, and fix $\eta$ = 0.1.
The results for approximating $M^{1/2}\bm{b}$ are summarized in Table 2 and Fig. 5, and the general trends align with the conclusions drawn form the 2D Laplacian example. However, the matrix in this case is no longer symmetric. From the plot of the error versus iteration count in Fig. 6, we see that as the condition number increases, the convergence rate closely follows a $k^{-3/4}$ relationship with respect to the iteration count. This observation is consistent with the theoretical result established in Theorem 3.8. Meanwhile, for a fixed $k=480$, Fig. 6 indicates that for large condition numbers, the observed trend closely follows $\kappa^{7/2}$.
Table 2: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for the convection-diffusion problem.
$n$ $\kappa(M)$ Relative Error Iteration Counts $k$
500 67510.90 3.99e-08 495
600 97087.72 3.31e-08 594
700 132022.80 2.85e-08 693
800 172316.14 2.53e-08 792
900 217967.73 2.29e-08 891
1000 268977.57 2.10e-08 990
1100 325345.68 1.94e-08 1089
1200 387072.03 1.81e-08 1188
1300 454156.65 1.71e-08 1287
Figure 5: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for the convection-diffusion problem.
Figure 6: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for the convection-diffusion problem.
@@
Figure 7: Log-log plot of relative error with $k=480$ against condition number for the convection-diffusion problem.
4.3 General matrices
Next, we consider positive definite but non-symmetric dense matrices. The matrix $M$ is constructed as follows:
(18) $M=M_{0}+\alpha I=USV^{T}+\alpha I,$
where $M_{0}$ is a low-rank matrix with rank $m$, $U$ and $V$ are orthonormal matrices generated randomly, and $S$ is an $m\times m$ diagonal matrix whose entries decay logarithmically from 1 to $\frac{1}{\beta}$, where $\beta$ is a given constant. A shift term, $\alpha I$, is introduced to ensure positive-definiteness of $M$. Specifically, the shift parameter $\alpha$ is defined as: $\alpha=-(1+5\times 10^{-3})\min\operatorname{Re}(\lambda_{\min}(M_{0})).$ As a result, the condition number of the $M$ is influenced by both $\alpha$ and $\beta$. In this test, we set the rank to $m=500$, and the size of the matrix $M$ is $5000\times 5000$. The parameter $\beta$ takes values of 400, 600, 800, 1000, 1200, 1400, 1600, 1800, 2000. As before, $\bm{b}$ is taken to be the vector of all ones.
In Table 3, we display the corresponding results. As the relative error remains around $10^{-5}$, the number of iterations generally increases with the condition number. Similarly, the normalized residual for solving $M\bm{x}=\bm{b}$ provides an effective stopping criterion for the Arnoldi iteration in the evaluation of $M^{1/2}\bm{b}$ (Fig. 8). As seen in Fig. 9, the error decay of the Arnoldi method closely follows the reference slope before the transition to superlinear convergence. However, we observe that with the iteration count fixed at $k=35$, the fitted slope is approximately $0.9$, which is much smaller than the theoretical value of $7/2$. These observations indicate that the theoretical results established in Theorem 3.13 yield a reasonably tight estimate for the iteration count $k$, but only a loose upper bound with respect to the condition number for certain type of matrices. This may be due to the fact that the bounds could depend not only on the condition number and iteration counts but also on intrinsic spectral characteristics or a specific structure of the matrices.
Table 3: Condition number of $M$, iteration counts, and relative error of the Arnoldi iteration for general matrices constructed via Eq. 18.
$\kappa(M)$ Relative Error Iteration Counts $k$
2192.38 2.40e-05 38
5375.35 2.90e-05 41
9325.42 2.62e-05 56
28826.11 3.86e-05 38
32437.66 2.45e-05 47
47395.74 2.87e-05 58
52845.93 6.31e-05 36
166131.16 5.27e-05 63
188372.52 1.03e-04 51
Figure 8: Semi-log plot of the normalized residual for FOM (*–) and the relative error for Arnoldi (o–) with different condition number against iteration count $k$ for general matrices constructed via Eq. 18.
@@ 4.3 General matrices
Figure 9: Log-log plot of relative error decay with varying condition numbers against iteration count $k$ for general matrices constructed via Eq. 18.
Figure 10: Log-log plot of relative error with $k=35$ against condition number for general matrices constructed via Eq. 18.
@@ 4.4 Practical application
The study of particulate suspensions is of great interest, and plays critical roles in diverse fields such as materials design [27], biological systems [27], micro-robotics [37], and food processing [21]. Immersed in a fluid environment, particles’ mobility is dictated by the hydrodynamic interactions between the particles. At micro scales, particles are also subject to thermal fluctuations, in the form of Brownian motion. The configuration of particles evolves over time and is governed by the overdamped limit of Langevin dynamics [14]. Thus, the computation of particles’ Brownian displacements crucially rely on the mobility matrix, and its square root, both dependent on all particles’ instantaneous positions. Efficiently computing these two matrices remains a significant computational challenge.
@@
Since hydrodynamic interactions depend on the separation distance between particles, the particle interactions weaken as the distance increases [25]. As a result, far-field blocks in the mobility matrix exhibit low-rank structure, and can then be efficiently approximated. To this end, the Hydrodynamic Interaction Graph Neural Network (HIGNN) framework, proposed in [25, 24], maps the many-body hydrodynamic interactions onto a graph neural network. By employing hierarchical-matrices using adaptive cross approximation [17], the computational complexity for using HIGNN to compute particles’ velocities given any applied forces can be reduced to nearly linear [26].
@@
In the following tests, $M$ is the mobility matrix obtained from the HIGNN model and has been preprocessed into a hierarchical-matrix form. Also, we assume that a uniform gravitational force $\bm{b}$ is applied to each particle. This numerical example focuses on evaluating the effectiveness of the Arnoldi iteration on computing $M^{1/2}$ in a long-time dynamic simulation of particles’ sedimentation. In the simulations, the explicit Euler method was adopted as the temporal integrator to update particles’ positions from their velocities, with a time step of $\Delta t=0.005$. At each time step, the HIGNN is applied to predict the velocities of all particles. To generate the initial cubic-lattice like configuration of particles, the particles are first positioned at grid points with a spacing of 4, and then each particle is randomly perturbed with a displacement of magnitude 0.2 from its original position. With this initial configuration, we simulate particles sedimenting in an unbounded domain. The number of particles, $N$, considered here varies from $20^{3}$ to $100^{3}$. For such large systems, direct computation of the square root of the mobility matrix $M$ can be computationally expensive. To address this and evaluate the accuracy of the Arnoldi iteration in this context, we apply the Arnoldi process twice to approximate $M\bm{b}=M^{1/2}(M^{1/2}\bm{b})$. Thus, the relative error is defined as $\frac{\|M\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\operatorname{Arn}_{k}(x^{1/2};M,\bm{b}))\|}{\|M\bm{b}\|}$.
@@
Figure 11: Snapshots of 8000 particles sedimenting in an unbounded domain at times $t=0.005$, $0.25$, and $1$, respectively, simulated using the HIGNN. The color is correlated to the particles’ velocity magnitude.
Table 4: Arnoldi Approximation for lattice structure system with constant particle concentration.
$N$ Iteration Counts $k$ Relative Error
@@
The snapshots at the 1st, 100th, and 200th time steps from the simulation of a $20^{3}$ particle system are presented in Fig. 11. Letting $M$ be the mobility matrix at the 100th time step, Table 4 records the number of iterations $k$ required to compute $M^{1/2}\bm{b}$ and the relative error as the number of particles $N$ increases from $20^{3}$ to $100^{3}$. Note that in this scenario, since the initial spacing between particles is kept the same as $N$ increases, a consistent particle concentration is maintained. One sees that as the number of particles increases, the number of iterations required to compute the matrix square root applied to a vector remains relatively low and does not grow significantly, while maintaining a relative error on the order of $10^{-4}$. This is because the matrix is well conditioned. Furthermore, we depict the relationship between the number of particles and the runtime on a log-log scale in Fig. 12. The plot implies that the Arnoldi approximation, based on the hierarchical matrix representation, exhibits an $O(N\log N)$ time complexity.
Figure 12: Log-log plot of runtime against number of particles for lattice structure system.
Next, we consider $40^{3}$ particles arranged in a similar initial cubic lattice as above, but with the grid spacing $d$ varying from 3 to 5 (still perturbed by a displacement of 0.2). With this change, we consider different particle concentrations; when $d$ increases, the particle concentration decreases, thus lowering the required rank of the approximation to the corresponding blocks. As shown in Table 5, lower particle concentration leads to reduced runtime for the Arnoldi approximation. However, the required number of iterations remains relatively small and stable compared to the matrix size. In this test, $M$ corresponds to the mobility matrix attained at the 50th time step.
Table 5: Arnoldi Approximation for lattice structure with different particle concentration.
@@ 5 Concluding remarks
In this work, we derive a priori error estimates of the Arnoldi process for approximating the matrix square root applied to a vector, extending the results in [6] for Hermitian matrices to non-Hermitian cases. Moreover, we provide an error bound based on a data-sparse representation of the matrix. Through numerical experiments, we investigate the convergence behavior of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ across different matrix types. The observations indicate that the theoretical results presented in Theorem 3.8 appear to be sharp in terms of iteration count, yet remain a loose upper bound with respect to the condition number for certain classes of matrices. Additionally, we consider the mobility matrix arising from particulate suspension simulations, where the matrix is approximated into a hierarchical matrix. The numerical results demonstrate the effectiveness of our approach for large-scale problems. Future extensions of this paper include deriving a sharper dependence on the condition number for approximating the matrix square root applied to a vector via Arnoldi iteration, as well as investigating the error behavior of GMRES in the approximation of $f(M)\bm{b}$ for general functions $f$.
@@ Acknowledgments
The authors would like to thank Zhan Ma and Zisheng Ye for their valuable discussions and assistance with the code implementation. The authors also gratefully acknowledge support from the National Science Foundation under Grant No. DMS-2208267 and the Army Research Office under Grant No. W911NF2310256.

### v3

@@ 4.2 Scaling behavior
where $u_{i}$ represents the numerical approximation to $u(x_{i})$. Let $M$ be the tridiagonal matrix obtained from the above finite-difference discretization of $L$, which is non-Hermitian.
Here, we take $\bm{b}$ to be the vector with all entries equal to one, and choose $\eta$ = 0.1. In this case, $M$ exhibits a relatively dispersed eigenvalue distribution. We compare different error bounds for approximating the square root of $M$ using the Arnoldi process described above (see Fig. 10). Similar phenomena as in Section 4.1 can also be seen here. Although the bound Eq. 11 is loose, its overall trend matches that of the true error, providing insights into the inherent scaling behavior.
@@
Figure 10: Error of the Arnoldi process for a convection-diffusion problem with $n=500$.
Next, we examine how the error depends on $\sigma_{\max}(M)$ and $k$, and whether the observed behavior follows the scaling predicted by bound (11). To this end, we introduce the scaling term derived from (11), defined as
$\frac{2^{1/4}\pi}{\Gamma\!(3/4)}\frac{2k-3}{2k}\frac{\left\|M^{1/2}\bm{b}-\operatorname{Arn}_{k}(f;M,\bm{b})\right\|}{\left\|\bm{\xi}_{0}^{k}\right\|}.$
In the following tests, we use the stopping criterion in Remark 3.10 with tol = 0.05.
The results of approximating $M^{1/2}\bm{b}$ for matrices of varying dimensions arising from the convection-diffusion problem are summarized in Table 1. We see that as $n$ increases, maintaining the same accuracy requires a larger number of iterations. The log-log plot of the scaling term decaying against the iteration count $k$ is shown in Fig. 11. The convergence rate exhibits a decay rate of approximately $\mathcal{O}(k^{-3/4})$ with respect to the iteration count $k$. This observation is consistent with the theoretical result established in Theorem 3.11. Additionally, we fix $k=300,500,700$ and plot the scaling term against the largest singular value on a log-log scale in Fig. 12. The results show that the fitted slopes vary with $k$, but, as expected, they all fall below the theoretical prediction.
Table 1: Condition number of $M$ and iteration counts $n$ and errors of the Arnoldi process for a convection-diffusion problem.
$n$ $\sigma_{\max}(M)$ error Iteration Counts $k$
1000 202320.64 0.03083 889
1200 291138.58 0.03053 1071
1400 396074.49 0.03061 1253
1600 517128.36 0.03090 1435
1800 654300.20 0.03132 1617
2000 807590.00 0.03132 1800
@@
Figure 11: Log-log plot of error decay with varying condition numbers against iteration count $k$ for a convection-diffusion problem.
Figure 12: Log-log plot of the error with $k=150$ against the largest singular value for a convection-diffusion problem.
@@ 4.3 Practical application
The study of particulate suspensions is of great interest, and plays critical roles in diverse fields such as materials design [makey2020universality], biological systems [makey2020universality], micro-robotics [yang2020reconfigurable], and food processing [kim2020modeling]. Immersed in a fluid environment, particles’ mobility is dictated by the hydrodynamic interactions between the particles. At micro scales, particles are also subject to thermal fluctuations, in the form of Brownian motion. The configuration of particles evolves over time and is governed by the overdamped limit of Langevin dynamics [ermak1978brownian]. Thus, the computation of particles’ Brownian displacements crucially rely on the mobility matrix, and its square root, both dependent on all particles’ instantaneous positions. Efficiently computing these two matrices remains a significant computational challenge.
@@
Since hydrodynamic interactions depend on the separation distance between particles, the particle interactions weaken as the distance increases [ma2022fast]. As a result, far-field blocks in the mobility matrix exhibit low-rank structure, and can then be efficiently approximated. To this end, the Hydrodynamic Interaction Graph Neural Network (HIGNN) framework, proposed in [ma2022fast, ma2024shape], maps the many-body hydrodynamic interactions onto a graph neural network. By employing hierarchical-matrices via adaptive cross approximation [hackbusch2015hierarchical], the computational complexity for using HIGNN to compute particles’ velocities given any forces can be reduced to nearly linear [ma2025mathcal].
@@
In the following tests, $M$ is the mobility matrix obtained from the HIGNN model and has been preprocessed into a hierarchical-matrix form. Also, we assume that a uniform gravitational force $\bm{b}$ is applied to each particle. This numerical example focuses on evaluating the effectiveness of the Arnoldi iteration on computing $M^{1/2}$ in a long-time dynamic simulation of particles’ sedimentation. In the simulations, the explicit Euler method was adopted as the temporal integrator to update particles’ positions from their velocities, with a time step of $\Delta t=0.005$. At each time step, the HIGNN is applied to predict the velocities of all particles. To generate the initial cubic-lattice like configuration of particles, the particles are first positioned at grid points with a spacing of 4, and then each particle is randomly perturbed with a displacement of magnitude 0.2 from its original position. With this initial configuration, we simulate particles sedimenting in an unbounded domain. The number of particles, $N$, considered here varies from $20^{3}$ to $100^{3}$. For such large systems, direct computation of the square root of the mobility matrix $M$ can be computationally prohibitive. To address this challenge and also validate the accuracy of the Arnoldi approximation in this practical setting, we apply the Arnoldi process twice to approximate $M\bm{b}=M^{1/2}(M^{1/2}\bm{b})$. Thus, the relative error is defined as $\frac{\|M\bm{b}-\operatorname{Arn}_{k}(x^{1/2};M,\operatorname{Arn}_{k}(x^{1/2};M,\bm{b}))\|}{\|M\bm{b}\|}$.
@@
Figure 13: Snapshots of 8000 particles sedimenting in an unbounded domain at times $t=0.005$, $0.25$, and $1$, respectively, simulated using the HIGNN [ma2025mathcal]. The color is correlated to the particles’ velocity magnitude.
Table 2: Iteration counts and errors of the Arnoldi process for a particulate suspension problem with varying particle number NN at fixed concentration.
NN Iteration Counts $k$ Relative Error
@@
The snapshots at the 1st, 100th, and 200th time steps from the simulation of a $20^{3}$ particle system are presented in Fig. 13. Letting $M$ be the mobility matrix at the 100th time step, Table 2 records the number of iterations $k$ required to compute $M^{1/2}\bm{b}$ and the relative error as the number of particles $N$ increases from $20^{3}$ to $100^{3}$. Note that in this scenario, since the initial spacing between particles is kept the same as $N$ increases, a fixed particle concentration is maintained. One sees that as the number of particles increases, the number of iterations required to compute the matrix square root applied to a vector remains relatively low and does not grow significantly, while maintaining a relative error on the order of $10^{-4}$. This is because the matrix is well conditioned. Furthermore, we depict the relationship between the number of particles and the runtime on a log-log scale in Fig. 14. The plot implies that the Arnoldi approximation, based on the hierarchical matrix representation, exhibits an $O(N\log N)$ time complexity.
Figure 14: Log-log plot of runtime versus number of particles NN at fixed particle concentration.
Next, we consider $40^{3}$ particles arranged in a similar initial cubic lattice as above, but with the grid spacing $d$ varying from 3 to 5 (still perturbed by a displacement of 0.2). With this change, we can examine varying particle concentrations; as $d$ increases, the particle concentration decreases, thus lowering the required rank of the approximation to the corresponding blocks. As shown in Table 3, lower particle concentration leads to reduced runtime for the Arnoldi approximation. Also, the required number of iterations remains relatively small and stable compared to the matrix size. In this test, $M$ corresponds to the mobility matrix attained at the 50th time step.
Table 3: Iteration counts, errors, and runtime of the Arnoldi process for a particulate suspension problem with varying particle concentration at $N=40^{3}$.
@@ 5 Concluding remarks
In this work, we derive error estimates of the Arnoldi process for approximating the matrix square root applied to a vector, extending the results reported in [chen2022error] for Hermitian matrices to non-Hermitian cases. In addition, we also provide a sharper bound for the Hermitian case, improving the state-of-the-art result. Moreover, we provide an error bound based on a perturbed matrix. Through a series of numerical experiments, we systematically assess the convergence behaviors of the Arnoldi iteration for computing $M^{1/2}\bm{b}$ across different matrix types. The findings indicate that the theoretical results presented in Theorem 3.8 appear to be sharp in terms of iteration count, yet remain a loose upper bound with respect to the largest singular value for certain classes of matrices. Furthermore, we consider the mobility matrix arising from particulate suspension simulations, where the matrix is approximated into a hierarchical form. The numerical results demonstrate the effectiveness of our approach for large-scale problems. Future extensions of this work include deriving a sharper dependence on the largest singular value for approximating the matrix square root applied to a vector via Arnoldi iteration, as well as investigating the error behaviors of GMRES in the approximation of $f(M)\bm{b}$ for general functions $f$.
@@ Acknowledgments
The authors would like to thank Zhan Ma and Zisheng Ye for their valuable discussions and assistance with the code implementation. The authors also gratefully acknowledge funding support from the National Science Foundation under Grant No. DMS-2208267 and the Army Research Office under Grant No. W911NF2310256.